# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v1

@@ 8.2. Trends and Opportunities
Based on the deployable NSVAD partitioned architecture depicted in Fig. 12, we have summarized the trends and open opportunities from hardware, system, algorithm, and application layers by incorporating the development trends of artificial intelligence and communication technologies, as well as the application requirements of NSVAD in smart cities and mobile internet, aiming to guide researchers in various fields to engage in relevant work.
@@ Acknowledgements.
This work has been supported in part by the China Mobile Research Fund of the Chinese Ministry of Education under Grant KEH2310029, the National Natural Science Foundation of China under Grant 62250410368, and the Specific Research Fund of the Innovation Platform for Academicians of Hainan Province under Grant YSPTZX202314. Additionally, this work has received backing from the Shanghai Key Research Laboratory of NSAI and the Joint Laboratory on Networked AI Edge Computing Fudan University-Changan. The work of Yang Liu is financially supported in part by the China Scholarship Council (File No. 202306100221). The authors would like to express their gratitude to the anonymous reviewers for their insightful comments and valuable suggestions. We appreciate Hao Yang, Liangyu Teng, and Yuntian Shi for their invaluable assistance and thorough review of this article.References
@@
- Adam et al. (2008) Amit Adam, Ehud Rivlin, Ilan Shimshoni, and Daviv Reinitz. 2008. Robust real-time unusual event detection using multiple fixed-location monitors. IEEE transactions on pattern analysis and machine intelligence 30, 3 (2008), 555–560.
- Aich et al. (2023) Abhishek Aich, Kuan-Chuan Peng, and Amit K Roy-Chowdhury. 2023. Cross-domain video anomaly detection without target domain adaptation. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 2579–2591.
@@
- Cao et al. (2023) Congqi Cao, Yue Lu, Peng Wang, and Yanning Zhang. 2023. A New Comprehensive Benchmark for Semi-supervised Video Anomaly Detection and Anticipation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 20392–20401.
@@
- Cheng et al. (2023) Kai Cheng, Yang Liu, and Xinhua Zeng. 2023. Learning graph enhanced spatial-temporal coherence for video anomaly detection. IEEE Signal Processing Letters 30 (2023), 314–318.
@@
- Cui et al. (2011) Xinyi Cui, Qingshan Liu, Mingchen Gao, and Dimitris N Metaxas. 2011. Abnormal detection using interaction energy potentials. In CVPR 2011. IEEE, 3161–3167.
@@
- Li et al. (2024) Chaobo Li, Hongjun Li, and Guoan Zhang. 2024. Cross-modality integration framework with prediction, perception and discrimination for video anomaly detection. Neural Networks 172 (2024), 106138.
@@
- Liu et al. (2024a) Yang Liu, Bobo Ju, Dingkang Yang, Liyuan Peng, Di Li, Peng Sun, Chengfang Li, Hao Yang, Jing Liu, and Liang Song. 2024a. Memory-enhanced spatial-temporal encoding framework for industrial anomaly detection system. Expert Systems with Applications (2024), 123718. https://doi.org/10.1016/j.eswa.2024.123718
@@
- Liu et al. (2022b) Yang Liu, Jing Liu, Wei Ni, and Liang Song. 2022b. Abnormal Event Detection with Self-guiding Multi-instance Ranking Framework. In 2022 International Joint Conference on Neural Networks (IJCNN). IEEE, 01–07.
@@
- Liu et al. (2023c) Yang Liu, Zhaoyang Xia, Mengyang Zhao, Donglai Wei, Yuzheng Wang, Liu Siao, Bobo Ju, Gaoyun Fang, Jing Liu, and Liang Song. 2023c. Learning Causality-inspired Representation Consistency for Video Anomaly Detection. In Proceedings of the 31st ACM International Conference on Multimedia. 203–212.
- Liu et al. (2024c) Yang Liu, Dingkang Yang, Yan Wang, Jing Liu, Jun Liu, Azzedine Boukerche, Peng Sun, and Liang Song. 2024c. Generalized video anomaly event detection: Systematic taxonomy and comparison of deep models. Comput. Surveys (2024).
@@
- Liu et al. (2023b) Zuhao Liu, Xiao-Ming Wu, Dian Zheng, Kun-Yu Lin, and Wei-Shi Zheng. 2023b. Generating anomalies for video anomaly detection with prompt-based feature mapping. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 24500–24510.
@@
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In International conference on machine learning. PMLR, 8748–8763.
@@
- Ramachandra and Jones (2020) Bharathkumar Ramachandra and Michael Jones. 2020. Street Scene: A new dataset and evaluation protocol for video anomaly detection. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 2569–2578.
@@
- Ravanbakhsh et al. (2018) Mahdyar Ravanbakhsh, Moin Nabi, Hossein Mousavi, Enver Sangineto, and Nicu Sebe. 2018. Plug-and-play cnn for crowd motion analysis: An application in abnormal event detection. In 2018 IEEE Winter Conference on Applications of Computer Vision (WACV). IEEE, 1689–1698.
- Ravanbakhsh et al. (2017) Mahdyar Ravanbakhsh, Moin Nabi, Enver Sangineto, Lucio Marcenaro, Carlo Regazzoni, and Nicu Sebe. 2017. Abnormal event detection in videos using generative adversarial nets. In 2017 IEEE international conference on image processing (ICIP). IEEE, 1577–1581.
@@
- Tran and Hogg (2017) Hanh TM Tran and David Hogg. 2017. Anomaly detection using a convolutional winner-take-all autoencoder. In Proceedings of the British Machine Vision Conference 2017. British Machine Vision Association.
- Turchini et al. (2017) Francesco Turchini, Lorenzo Seidenari, and Alberto Del Bimbo. 2017. Convex polytope ensembles for spatio-temporal anomaly detection. In International Conference on Image Analysis and Processing. Springer, 174–184.
@@
- Wang et al. (2021) Xuanzhao Wang, Zhengping Che, Bo Jiang, Ning Xiao, Ke Yang, Jian Tang, Jieping Ye, Jingyu Wang, and Qi Qi. 2021. Robust unsupervised video anomaly detection by multipath frame prediction. IEEE transactions on neural networks and learning systems (2021).
@@
- Wu et al. (2024a) Peng Wu, Jing Liu, Xiangteng He, Yuxin Peng, Peng Wang, and Yanning Zhang. 2024a. Toward Video Anomaly Retrieval From Video Anomaly Detection: New Benchmarks and Model. IEEE Transactions on Image Processing 33 (2024), 2213–2225.
@@
- Wu et al. (2024b) Peng Wu, Xuerong Zhou, Guansong Pang, Yujia Sun, Jing Liu, Peng Wang, and Yanning Zhang. 2024b. Open-Vocabulary Video Anomaly Detection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
- Wu et al. (2024c) Peng Wu, Xuerong Zhou, Guansong Pang, Lingru Zhou, Qingsen Yan, Peng Wang, and Yanning Zhang. 2024c. Vadclip: Adapting vision-language models for weakly supervised video anomaly detection. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 6074–6082.
@@
- Yang et al. (2024) Zhiwei Yang, Jing Liu, and Peng Wu. 2024. Text Prompt with Normality Guidance for Weakly Supervised Video Anomaly Detection. arXiv preprint arXiv:2404.08531 (2024).

### v2

@@ 8.2. Trends and Opportunities
Based on the sequential steps of NSVAD depicted in Fig. 1, we have summarized the trends and open opportunities from hardware, system, algorithm, and application layers by incorporating the development trends of artificial intelligence and communication technologies, as well as the application requirements of NSVAD in smart cities and mobile internet, aiming to guide researchers in various fields to engage in relevant work.
@@ Acknowledgements.
This work is supported in part by the China Mobile Research Fund of MOE under Grant KEH2310029, NSFC under Grant 62250410368, and China Scholarship Council (File No. 202306100221). The authors would like to express their gratitude to the anonymous reviewers for their insightful comments and valuable suggestions.References
@@
- Al-Lahham et al. (2024) Anas Al-Lahham, Muhammad Zaigham Zaheer, Nurbek Tastan, and Karthik Nandakumar. 2024. Collaborative Learning of Anomalies with Privacy (CLAP) for Unsupervised Video Anomaly Detection: A New Baseline. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 12416–12425.
@@
- Chen et al. (2023) Weiling Chen, Keng Teck Ma, Zi Jian Yew, Minhoe Hur, and David Aik-Aun Khoo. 2023. TEVAD: Improved video anomaly detection with captions. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 5548–5558.
- Cheng et al. (2020) Hang Cheng, Ximeng Liu, Huaxiong Wang, Yan Fang, Meiqing Wang, and Xiaopeng Zhao. 2020. SecureAD: A secure video anomaly detection framework on convolutional neural network in edge computing environment. IEEE Transactions on Cloud Computing 10, 2 (2020), 1413–1427.
- Cheng et al. (2023a) Kai Cheng, Yang Liu, and Xinhua Zeng. 2023a. Learning graph enhanced spatial-temporal coherence for video anomaly detection. IEEE Signal Processing Letters 30 (2023), 314–318.
- Cheng et al. (2024a) Kai Cheng, Yaning Pan, Yang Liu, Xinhua Zeng, and Rui Feng. 2024a. Denoising Diffusion-Augmented Hybrid Video Anomaly Detection via Reconstructing Noised Frames. In Proceedings of the Thirty-Third International Joint Conference on Artificial Intelligence, IJCAI-24. 695–703.
- Cheng et al. (2024b) Kai Cheng, Xinhua Zeng, Yang Liu, Yaning Pan, and Xinzhe Li. 2024b. Normality learning reinforcement for anomaly detection in surveillance videos. Knowledge-Based Systems 297 (2024), 111942.
- Cheng et al. (2023b) Kai Cheng, Xinhua Zeng, Yang Liu, Mengyang Zhao, Chengxin Pang, and Xing Hu. 2023b. Spatial-temporal graph convolutional network boosted flow-frame prediction for video anomaly detection. In ICASSP 2023-2023 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 1–5.
@@
- Dong et al. (2024) Shi Dong, Junxiao Tang, Khushnood Abbas, Ruizhe Hou, Joarder Kamruzzaman, Leszek Rutkowski, and Rajkumar Buyya. 2024. Task offloading strategies for mobile edge computing: A survey. Computer Networks (2024), 110791.
@@
- Duong et al. (2023) Huu-Thanh Duong, Viet-Tuan Le, and Vinh Truong Hoang. 2023. Deep learning-based anomaly detection in video surveillance: A survey. Sensors 23, 11 (2023), 5024.
@@
- Fioresi et al. (2023) Joseph Fioresi, Ishan Rajendrakumar Dave, and Mubarak Shah. 2023. Ted-spad: Temporal distinctiveness for self-supervised privacy-preservation for video anomaly detection. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 13598–13609.
@@
- Gomes et al. (2023) Carlos Gomes, Roberto Azevedo, and Christopher Schroers. 2023. Video compression with entropy-constrained neural representations. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 18497–18506.
@@
- Huang et al. (2024a) Chao Huang, Chengliang Liu, Jie Wen, Lian Wu, Yong Xu, Qiuping Jiang, and Yaowei Wang. 2024a. Weakly Supervised Video Anomaly Detection via Self-Guided Temporal Discriminative Transformer. IEEE Transactions on Cybernetics (2024).
- Huang et al. (2022) Chao Huang, Yabo Liu, Zheng Zhang, Chengliang Liu, Jie Wen, Yong Xu, and Yaowei Wang. 2022. Hierarchical Graph Embedded Pose Regularity Learning via Spatio-Temporal Transformer for Abnormal Behavior Detection. In 30th ACM International Conference on Multimedia. 307–315.
- Huang et al. (2024b) Chao Huang, Jie Wen, Chengliang Liu, and Yabo Liu. 2024b. Long Short-Term Dynamic Prototype Alignment Learning for Video Anomaly Detection. Proceedings of the Thirty-Two International Joint Conference on Artificial Intelligence (IJCAI-24) (April 2024).
- Huang et al. (2023) Chao Huang, Jie Wen, Yong Xu, Qiuping Jiang, Jian Yang, Yaowei Wang, and David Zhang. 2023. Self-Supervised Attentive Generative Adversarial Networks for Video Anomaly Detection. IEEE Transactions on Neural Networks and Learning Systems 34, 11 (Nov. 2023), 9389–9403.
@@
- Karim et al. (2024) Hamza Karim, Keval Doshi, and Yasin Yilmaz. 2024. Real-time weakly supervised video anomaly detection. In Proceedings of the IEEE/CVF winter conference on applications of computer vision. 6848–6856.
- Kenton and Toutanova (2019) Jacob Devlin Ming-Wei Chang Kenton and Lee Kristina Toutanova. 2019. Bert: Pre-training of deep bidirectional transformers for language understanding. In Proceedings of naacL-HLT, Vol. 1. Minneapolis, Minnesota, 2.
- Kim et al. (2023) Jaehyun Kim, Seongwook Yoon, Taehyeon Choi, and Sanghoon Sull. 2023. Unsupervised video anomaly detection based on similarity with predefined text descriptions. Sensors 23, 14 (2023), 6256.
@@
- Li et al. (2024a) Chaobo Li, Hongjun Li, and Guoan Zhang. 2024a. Cross-modality integration framework with prediction, perception and discrimination for video anomaly detection. Neural Networks 172 (2024), 106138.
- Li et al. (2024b) Chengfang Li, Yang Liu, Jielin Li, Yuhao Miao, Jing Liu, and Liang Song. 2024b. Decoding Bilingual EEG Signals With Complex Semantics Using Adaptive Graph Attention Convolutional Network. IEEE Transactions on Neural Systems and Rehabilitation Engineering (2024).
- Li et al. (2024c) Chengfang Li, Hanqi Wang, Yang Liu, Xiaoguang Zhu, and Liang Song. 2024c. Silent EEG classification using cross-fusion adaptive graph convolution network for multilingual neurolinguistic signal decoding. Biomedical Signal Processing and Control 87 (2024), 105524.
@@
- Liu et al. (2023c) Jing Liu, Yang Liu, Di Li, Hanqi Wang, Xiaohong Huang, and Liang Song. 2023c. DSDCLA: Driving style detection via hybrid CNN-LSTM with multi-level attention fusion. Applied Intelligence 53, 16 (2023), 19237–19254.
- Liu et al. (2023d) Jing Liu, Yang Liu, Wei Zhu, Xiaoguang Zhu, and Liang Song. 2023d. Distributional and spatial-temporal robust representation learning for transportation activity recognition. Pattern Recognition 140 (2023), 109568.
- Liu et al. (2024c) Jing Liu, Donglai Wei, Yang Liu, Sipeng Zhang, Tong Yang, and Victor C. M. Leung. 2024c. SCMM: Calibrating Cross-modal Representations for Text-Based Person Search. arXiv:2304.02278
@@
- Liu et al. (2024a) Yang Liu, Bobo Ju, Dingkang Yang, Liyuan Peng, Di Li, Peng Sun, Chengfang Li, Hao Yang, Jing Liu, and Liang Song. 2024a. Memory-enhanced spatial-temporal encoding framework for industrial anomaly detection system. Expert Systems with Applications (2024), 123718.
@@
- Liu et al. (2023b) Yang Liu, Di Li, Wei Zhu, Dingkang Yang, Jing Liu, and Liang Song. 2023b. MSN-net: Multi-scale normality network for video anomaly detection. In ICASSP 2023-2023 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 1–5.
@@
- Liu et al. (2022b) Yang Liu, Jing Liu, Mengyang Zhao, Shuang Li, and Liang Song. 2022b. Collaborative Normality Learning Framework for Weakly Supervised Video Anomaly Detection. IEEE Transactions on Circuits and Systems II: Express Briefs 69, 5 (2022), 2508–2512.
@@
- Liu et al. (2022d) Yang Liu, Jing Liu, Xiaoguang Zhu, Donglai Wei, Xiaohong Huang, and Liang Song. 2022d. Learning Task-Specific Representation for Video Anomaly Detection with Spatial-Temporal Attention. In ICASSP 2022-2022 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 2190–2194.
- Liu et al. (2023e) Yang Liu, Zhaoyang Xia, Mengyang Zhao, Donglai Wei, Yuzheng Wang, Liu Siao, Bobo Ju, Gaoyun Fang, Jing Liu, and Liang Song. 2023e. Learning Causality-inspired Representation Consistency for Video Anomaly Detection. In 31st ACM International Conference on Multimedia. 203–212.
- Liu et al. (2023f) Yang Liu, Dingkang Yang, Gaoyun Fang, Yuzheng Wang, Donglai Wei, Mengyang Zhao, Kai Cheng, Jing Liu, and Liang Song. 2023f. Stochastic video normality network for abnormal event detection in surveillance videos. Knowledge-Based Systems 280 (2023), 110986.
- Liu et al. (2024d) Yang Liu, Dingkang Yang, Yan Wang, Jing Liu, Jun Liu, Azzedine Boukerche, Peng Sun, and Liang Song. 2024d. Generalized video anomaly event detection: Systematic taxonomy and comparison of deep models. Comput. Surveys 56, 7 (2024), 1–38.
@@
- Lv and Sun (2024) Hui Lv and Qianru Sun. 2024. Video anomaly detection and explanation via large language models. arXiv preprint arXiv:2401.05702 (2024).
@@
- Nayak et al. (2023) Rashmiranjan Nayak, Umesh Chandra Pati, and Santos Kumar Das. 2023. Video Anomaly Detection Using Self-Attention-Enabled Convolutional Spatiotemporal Autoencoder. In 2023 22nd International Symposium on Communications and Information Technologies (ISCIT). IEEE, 70–75.
- Nayak et al. (2024) Rashmiranjan Nayak, Umesh Chandra Pati, and Santos Kumar Das. 2024. A comprehensive review of datasets for detection and localization of video anomalies: a step towards data-centric artificial intelligence-based video anomaly detection. Multimedia Tools and Applications 83, 21 (2024), 59617–59674.
@@
- Noghre et al. (2024) Ghazal Alinezhad Noghre, Shanle Yao, Armin Danesh Pazho, Babak Rahimi Ardabili, Vinit Katariya, and Hamed Tabkhi. 2024. PHEVA: A Privacy-preserving Human-centric Video Anomaly Detection Dataset. arXiv preprint arXiv:2408.14329 (2024).
@@
- Pu and Wu (2022) Yujiang Pu and Xiaoyu Wu. 2022. Audio-Guided Attention Network for Weakly Supervised Violence Detection. In 2022 2nd International Conference on Consumer Electronics and Computer Engineering (ICCECE). IEEE, 219–223.
- Pu et al. (2024) Yujiang Pu, Xiaoyu Wu, Lulu Yang, and Shengjin Wang. 2024. Learning prompt-enhanced context features for weakly-supervised video anomaly detection. IEEE Transactions on Image Processing (2024).
@@
- Radford (2018) Alec Radford. 2018. Improving language understanding by generative pre-training. (2018).
- Radford et al. (2021a) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021a. Learning transferable visual models from natural language supervision. In International conference on machine learning. PMLR, 8748–8763.
- Radford et al. (2021b) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021b. Learning transferable visual models from natural language supervision. In International conference on machine learning. PMLR, 8748–8763.
@@
- Su et al. (2025) Yong Su, Yuyu Tan, Simin An, Meng Xing, and Zhiyong Feng. 2025. Semantic-driven dual consistency learning for weakly supervised video anomaly detection. Pattern Recognition 157 (2025), 110898.
- Su et al. (2023) Yong Su, Haohao Zhu, Yuyu Tan, Simin An, and Meng Xing. 2023. Prime: privacy-preserving video anomaly detection via motion exemplar guidance. Knowledge-Based Systems 278 (2023), 110872.
@@
- Tang et al. (2024) Jiaqi Tang, Hao Lu, Ruizheng Wu, Xiaogang Xu, Ke Ma, Cheng Fang, Bin Guo, Jiangbo Lu, Qifeng Chen, and Ying-Cong Chen. 2024. Hawk: Learning to Understand Open-World Video Anomalies. arXiv preprint arXiv:2405.16886 (2024).
- Teng et al. (2024) Liangyu Teng, Yang Liu, Jing Liu, and Liang Song. 2024. End-Cloud Collaboration Framework for Advanced AI Customer Service in E-commerce. arXiv preprint arXiv:2410.07122 (2024).
@@
- Wang et al. (2021b) Hanqi Wang, Xing Hu, Liang Song, Guanhua Zhang, Yang Liu, Jing Liu, and Linhua Jiang. 2021b. Stack Multiple Shallow Autoencoders into a Strong One: A New Reconstruction-Based Method to Detect Anomaly. In International Conference on Neural Information Processing. Springer, 103–115.
@@
- Wang et al. (2021a) Xuanzhao Wang, Zhengping Che, Bo Jiang, Ning Xiao, Ke Yang, Jian Tang, Jieping Ye, Jingyu Wang, and Qi Qi. 2021a. Robust unsupervised video anomaly detection by multipath frame prediction. IEEE transactions on neural networks and learning systems (2021).
- Wang et al. (2023) Yang Wang, Tianying Liu, Jiaogen Zhou, and Jihong Guan. 2023. Video anomaly detection based on spatio-temporal relationships among objects. Neurocomputing 532 (2023), 141–151.
@@
- Wang et al. (2024a) Yan Wang, Shaoqi Yan, Yang Liu, Wei Song, Jing Liu, Yang Chang, Xinji Mai, Xiping Hu, Wenqiang Zhang, and Zhongxue Gan. 2024a. A Survey on Facial Expression Recognition of Static and Dynamic Emotions. arXiv preprint arXiv:2408.15777 (2024).
- Wang et al. (2024b) Yuzheng Wang, Dingkang Yang, Zhaoyu Chen, Yang Liu, Siao Liu, Wenqiang Zhang, Lihua Zhang, and Lizhe Qi. 2024b. De-confounded Data-free Knowledge Distillation for Handling Distribution Shifts. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 12615–12625.
- Wei et al. (2022b) Donglai Wei, Yang Liu, Xiaoguang Zhu, Jing Liu, and Xinhua Zeng. 2022b. MSAF: Multimodal Supervise-Attention Enhanced Fusion for Video Anomaly Detection. IEEE Signal Processing Letters 29 (2022), 2178–2182.
- Wei et al. (2022a) Dong-Lai Wei, Chen-Geng Liu, Yang Liu, Jing Liu, Xiao-Guang Zhu, and Xin-Hua Zeng. 2022a. Look, Listen and Pay More Attention: Fusing Multi-Modal Information for Video Violence Detection. In ICASSP 2022-2022 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 1980–1984.
- Wu et al. (2024a) Lian Wu, Chao Huang, Lunke Fei, Shuping Zhao, Jianchuan Zhao, Zhongwei Cui, and Yong Xu. 2024a. Video-Based Fall Detection Using Human Pose and Constrained Generative Adversarial Network. IEEE Transactions on Circuits and Systems for Video Technology 34, 4 (April 2024), 2179–2194.
- Wu et al. (2024b) Peng Wu, Jing Liu, Xiangteng He, Yuxin Peng, Peng Wang, and Yanning Zhang. 2024b. Toward Video Anomaly Retrieval From Video Anomaly Detection: New Benchmarks and Model. IEEE Transactions on Image Processing 33 (2024), 2213–2225.
@@
- Wu et al. (2024c) Peng Wu, Xuerong Zhou, Guansong Pang, Yujia Sun, Jing Liu, Peng Wang, and Yanning Zhang. 2024c. Open-Vocabulary Video Anomaly Detection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
- Wu et al. (2024d) Peng Wu, Xuerong Zhou, Guansong Pang, Zhiwei Yang, Qingsen Yan, Peng Wang, and Yanning Zhang. 2024d. Weakly supervised video anomaly detection and localization with spatio-temporal prompts. arXiv preprint arXiv:2408.05905 (2024).
- Wu et al. (2024e) Peng Wu, Xuerong Zhou, Guansong Pang, Lingru Zhou, Qingsen Yan, Peng Wang, and Yanning Zhang. 2024e. Vadclip: Adapting vision-language models for weakly supervised video anomaly detection. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 6074–6082.
@@
- Yan et al. (2023) Cheng Yan, Shiyu Zhang, Yang Liu, Guansong Pang, and Wenjun Wang. 2023. Feature prediction diffusion model for video anomaly detection. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 5527–5537.
@@
- Yang et al. (2024a) Dingkang Yang, Mingcheng Li, Dongling Xiao, Yang Liu, Kun Yang, Zhaoyu Chen, Yuzheng Wang, Peng Zhai, Ke Li, and Lihua Zhang. 2024a. Towards multimodal sentiment analysis debiasing via bias purification. arXiv preprint arXiv:2403.05023 (2024).
- Yang et al. (2024c) Hao Yang, Hongyuan Lu, Xinhua Zeng, Yang Liu, Xiang Zhang, Haoran Yang, Yumeng Zhang, Yiran Wei, and Wai Lam. 2024c. Stephanie: Step-by-Step Dialogues for Mimicking Human Interactions in Social Conversations. arXiv preprint arXiv:2407.04093 (2024).
@@
- Yang et al. (2023b) Kun Yang, Dingkang Yang, Jingyu Zhang, Mingcheng Li, Yang Liu, Jing Liu, Hanqi Wang, Peng Sun, and Liang Song. 2023b. Spatio-temporal domain awareness for multi-agent collaborative perception. In Proceedings of the IEEE/CVF International Conference on Computer Vision, 23383-23392.
- Yang et al. (2024b) Zhiwei Yang, Jing Liu, and Peng Wu. 2024b. Text Prompt with Normality Guidance for Weakly Supervised Video Anomaly Detection. arXiv preprint arXiv:2404.08531 (2024).
- Yang et al. (2023a) Zhiwei Yang, Jing Liu, Zhaoyang Wu, Peng Wu, and Xiaotao Liu. 2023a. Video event restoration based on keyframes for video anomaly detection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 14592–14601.
@@
- Yu et al. (2022) Jiashuo Yu, Jinyu Liu, Ying Cheng, Rui Feng, and Yuejie Zhang. 2022. Modality-Aware Contrastive Instance Learning with Self-Distillation for Weakly-Supervised Audio-Visual Violence Detection. In Proceedings of the 30th ACM International Conference on Multimedia. 6278–6287.
@@
- Zanella et al. (2024) Luca Zanella, Willi Menapace, Massimiliano Mancini, Yiming Wang, and Elisa Ricci. 2024. Harnessing Large Language Models for Training-free Video Anomaly Detection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 18527–18536.
- Zhang et al. (2024b) Huaxin Zhang, Xiang Wang, Xiaohao Xu, Xiaonan Huang, Chuchu Han, Yuehuan Wang, Changxin Gao, Shanjun Zhang, and Nong Sang. 2024b. GlanceVAD: Exploring Glance Supervision for Label-efficient Video Anomaly Detection. arXiv preprint arXiv:2403.06154 (2024).
- Zhang et al. (2024c) Huaxin Zhang, Xiaohao Xu, Xiang Wang, Jialong Zuo, Chuchu Han, Xiaonan Huang, Changxin Gao, Yuehuan Wang, and Nong Sang. 2024c. Holmes-VAD: Towards Unbiased and Explainable Video Anomaly Detection via Multi-modal LLM. arXiv preprint arXiv:2406.12235 (2024).
- Zhang et al. (2024a) Menghao Zhang, Jingyu Wang, Qi Qi, Haifeng Sun, Zirui Zhuang, Pengfei Ren, Ruilong Ma, and Jianxin Liao. 2024a. Multi-Scale Video Anomaly Detection by Multi-Grained Spatio-Temporal Representation Learning. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 17385–17394.
@@
- Zhao et al. (2024) Mengyang Zhao, Xinhua Zeng, Yang Liu, Jing Liu, and Chengxin Pang. 2024. Rethinking prediction-based video anomaly detection from local-global normality perspective. Expert Systems with Applications (2024), 125581.
@@
- Zhou et al. (2024) Yixuan Zhou, Yi Qu, Xing Xu, Fumin Shen, Jingkuan Song, and Heng Tao Shen. 2024. Batchnorm-based weakly supervised video anomaly detection. IEEE Transactions on Circuits and Systems for Video Technology (2024).